OpenStudy (anonymous):

on a triangle the height is 3/8 and the length is 2 1/3 what is the area yard?

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

one half base times height

OpenStudy (anonymous):

i.e. \[\frac{1}{2}\times \frac{3}{8}\times 2\tfrac{1}{3}\]

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

are ya there

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

did you step away